Determine the genotype at the DT locus for a population of 73 individuals who have "osmosis" syndrome and observe the following.

Genotype DD DT TT

Number of individuals DD=14 DT=38 TT=21

Is this population at Hardy-Weinberg Equilibrium?

X^2 critical= 5.99, df=2, a=0.05
Answer #1

Answer:

DD 14 = 28 D alleles

DT 38 = 38 D & 38 T alleles

TT 21 = 42 T alleles

Total alleles = 146 alleles

Total D alleles = 28+38 = 66

Frequency of D allele = 66 / 146 = 0.45

Total d alleles = 42+38 = 80

Frequency of d allele = 80 / 146 = 0.55

Expected number of each genotype:

DD = 0.45 * 0.45 * 73 = 14.78

DT = 2 * 0.45 * 0.55 * 73= 36.14

TT = 0.55 * 0.55 * 73 = 22.08

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

DD

14

14.78

-0.78

0.61

0.0412

DT

38

36.14

1.86

3.46

0.0957

TT

21

22.08

-1.08

1.17

0.0528

Total

73

73

0.1897

X^2 value = 0.19

Degrees of freedom = number of categories – 1

Df = 3-1 = 2

Critical value = 5.99

The chi-square value of 0.19 is less than the critical value of 5.99. So we can accept the null hypothesis. Hence the population is in Hardy-Weinberg equilibrium.